it took 3 hours to row a boat 18 km against the current. The return trip with the current took 1 1/2 hours. find the speed of the plane in still air.
I know d = r x t, but I am stuck on how to get the still air. I know to make both equations = to 18, but help.

Answer by richwmiller(17219)

r*t=d
r*3=18
(r-x)*3=18
r-x=6
(r+x)*3/2=18
r+x=12
r-x=6
2r=18
r=9 in still water
r+x=12
x=3
current=3
I assume they really mean still water not still air and that they mean boat not plane since they left in a boat not plane.